OpenStudy (anonymous):

Solve for the variable p - 1 = - 54

OpenStudy (anonymous):

what you are not getting in this question???

OpenStudy (anonymous):

add 1

OpenStudy (uri):

p - 1 = - 54 p =54 +1 Add 1 and p =55 then.